OpenStudy (vishweshshrimali5):

A plane passes through the points (6,-4,3) and (0,4,-3). The sum of its intercepts on the coordinate axes is zero. Its distance from the origin is

OpenStudy (anonymous):

x/a + y/b + z/c =1 a + b + c = 0 plug in the two pts to get two more eqns 3 variables and 3 equations, find a b and c

OpenStudy (vishweshshrimali5):

please explain me the whole answer him1618 and please with every step explained

OpenStudy (anonymous):

x/a + y/b + z/c =1 this is the intercept form of a plane sum of intercepts is zero so a+b+c=0

OpenStudy (vishweshshrimali5):

Okkkk

OpenStudy (vishweshshrimali5):

Then ?

OpenStudy (anonymous):

now plug in (6,4,3) into ur plane 6/a + 4/b + 3/c = 1 plug in (0,4,-3) now 4/b -3/c = 1 3 variables and three eqns

OpenStudy (anonymous):

solve them

OpenStudy (vishweshshrimali5):

Let me see if I can solve them

OpenStudy (vishweshshrimali5):

Where is the third equation ?